# PerlDiff: Controllable Street View Synthesis Using Perspective-Layout Diffusion Models

### v1

@@ 3.2 Object Controllability via Perl-based Controlling Module
Figure 4: Qualitative visualization on NuScenes [1] demonstrating the effects of Text Cross-attention. From left to right: day, night, and rain scenarios synthesized by PerlDiff, highlighting its adaptability to different lighting and weather conditions. For more examples, see the Appendix D.
3.3 Discussion
In contrast to previous approaches for autonomous driving such as BEVControl [38], MagicDrive [6], DrivingDiffusion [12], and Panacea [36], which employ a basic cross-attention mechanism to integrate controlling condition information, our PerlDiff leverages geometric priors via Perl masking map. This approach directs the generation of each object with its respective control information during the training phase, effectively countering the common misalignment between the attention map and condition information that often results in compromised image controllability. For instance, the attention map of BEVControl [38], illustrated in Fig. 3, demonstrates disorganized patterns and lacks precision. Conversely, our PerlDiff markedly enhances the accuracy of generated images and the granularity of condition information at the object level by ensuring meticulous guidance within the attention map. For more qualitative results, please see the Appendix D.
4 Experiments
We assess PerlDiff’s ability to control quality across multiple benchmarks, including multi-view 3D object detection, BEV segmentation and monocular 3D object detection. Subsequently, we conduct ablation studies to ascertain the contribution of each component within our proposed methodology.
4.1 Datasets
NuScenes [1] dataset comprises 1,000 urban street scenes, traditionally segmented into 700 for training, 150 for validation, and 150 for testing. Each scene features six high-resolution images (900$\times$1600), which together provide a complete 360-degree panoramic view of the surroundings. Additionally, NuScenes includes comprehensive road maps of the driving environment, featuring details such as lane markings and obstacles. We extend the class and road type annotations similar to MagicDrive [6] and NuScenes, incorporating ten object classes and eight road types for map rendering. To address the resolution limitations of the U-Net architecture in Stable Diffusion [25], we adopt image resolutions of 256$\times$384 as in BEVFormer [14], and 256$\times$704 following BEVFusion [15].
KITTI [7] dataset contains 3,712 images for training and 3,769 images for validation. KITTI dataset has only one perspective image and does not have road map information. Given the varied image resolutions in KITTI (approximately 375$\times$1242), we pad them to 384$\times$1280 for generative learning.
4.2 Main Results
In this subsection, we assess our PerlDiff’s generative quality through the perception results of several pre-trained methods: BEVFormer [14], BEVFusion [15], and StreamPETR [34] for multi-view 3D detection; CVT [41] for BEV segmentation; all trained on the NuScenes set, and MonoFlex [40] for monocular 3D detection trained on the KITTI set. Additionally, we leverage our synthesized dataset to enhance the performance of various 3D detection models (i.e., BEVFormer and StreamPETR) on the NuScenes test set, validating the effectiveness of our PerlDiff.
Table 1: Comparison of the controllability of street view image generation on NuScenes [1] validation set. Our replication of BEVControl*, serving as the baseline, employs standard cross-attention mechanisms contrary to Perl-based cross-attention utilized in our PerlDiff. Outcomes demonstrating superior performance are highlighted in bold.
@@ 4.2 Main Results
Oracle BEVFormer 27.06 41.89 0.54 70.35 33.36
Oracle BEVFusion 35.54 41.20 0.56 70.46 35.86
MagicDrive [6] BEVFusion 16.20 12.30 23.32 61.05 27.01
BEVControl* BEVFusion 23.89 9.98 19.61 0.94 60.74 22.47
PerlDiff (Ours) BEVFusion 25.06 15.24 24.05 0.78 61.26 27.13
BEVGen [30] 25.54 50.20 5.89
BEVControl [38] BEVFormer 24.85 19.64 28.68 0.78 60.80 26.86
BEVControl* BEVFormer 23.89 16.48 28.08 0.88 60.74 22.47
PerlDiff (Ours) BEVFormer 25.06 25.10 36.24 0.72 61.26 27.13
Table 2: Controllability comparison on KITTI [7] validation set, showcasing vehicle mAP obtained by MonoFlex [40] using data generated by our PerlDiff and the baseline BEVControl*. “NuScenes $\rightarrow$ KITTI” denotes initial training on NuScenes train set followed by fine-tuning on KITTI train set.
@@
Controllable Generation on NuScenes. To evaluate the effectiveness of our PerlDiff, we trained the model on NuScenes train set and subsequently generated a synthetic validation set using the provided road maps and 3D annotations. The controllability of PerlDiff was examined by deploying perception models, originally trained on the actual train set, onto our synthetic validation set. As summarized in Tab. 1, our PerlDiff outperforms competing methods on most metrics, as tested with BEVFormer [14] and BEVFusion [15]. In a rigorous comparison, we replicated BEVControl [38] using identical settings, with the exception of our innovative element: the Perl-based cross-attention mechanism. Our PerlDiff demonstrates notable improvements, with increments of 8.62%, 8.16% and 0.16% in the mean Average Precision (mAP), NuScenes detection score (NDS) and mean average orientation error (mAOE), respectively, compared to BEVControl* when using BEVFormer. With BEVFusion, it achieves gains of 5.26%, 4.44% and 0.16% in these metrics against BEVControl*, confirming its effectiveness at a resolution of $256\times 384$. The superiority of our PerlDiff is further affirmed by BEV segmentation [41] metrics, significantly outperforming BEVControl* with a 0.52% increase in Road mIoU and a 4.66% increase in Vehicle mIoU, validating the potency of the Perl-based cross-attention mechanism in improving the controllability of the scene.
Compared to the state-of-the-art MagicDrive [6], our method lags only in the FID (Fréchet Inception Distance) metric [8]. Despite this, prioritizing definitive control over generative images is crucial to their practical application in the real world. For perception models, excessive similarity in synthetic output can lead to data redundancy. our PerlDiff skillfully ensures a balance within the synthesized data spectrum, as reflected by its FID scores comparable to BEVControl*, positioning our framework as a prime choice for enhancing conditional diffusion models.
Controllable Generation on KITTI. The scarcity of training data in KITTI [7] often hampers a generative model’s capacity to discern the association between control information and the synthesis of images. To mitigate this limitation, we employ two distinct strategies for generating images within the KITTI framework: one strategy is direct training utilizing the KITTI train set, and the other involves preliminary training on the NuScenes train set with subsequent fine-tuning on the KITTI train set. In Tab. 2, we present the results of monocular 3D object detection on KITTI validation set, utilizing a pretrained MonoFlex [40] detector. In particular, BEVControl* shows substantially reduced detection efficacy with few training samples. In contrast, our PerlDiff maintains a consistently high level of data annotation performance, outperforming BEVControl* by a considerable margin. Thus, our model exhibits pronounced superiority in scenarios involving datasets with limited sizes. The visual comparison between BEVControl* and our PerlDiff is presented in Fig. 5. More visual results on KITTI can be found in the Appendix D.
Table 3: Performance comparison for the boosting performance of 3D detection models using synthesized dataset on NuScenes [1] test set using BEVFormer [14] and StreamPETR [34]. The “train + Real val” configuration serves as a benchmark, representing the ideal upper performance limit achievable. “Syn. val” represents the synthetic validation set generated by BEVControl. The numbers in parentheses indicate the performance disparity relative to the “train + Real val” configuration.
@@
Figure 5: Qualitative visualization comparison on KITTI [7]. Red markers denote instances where BEVControl* inaccurately generates output compared to PerlDiff and ground truth.
Boosting Perception Models Using Synthesized Dataset. Generative models have become widely acknowledged as effective tools for data augmentation, thereby improving the generalization capabilities of perception models. To evaluate this approach, we leverage our synthesized dataset to improve the performance of various detection models on NuScenes test set. The gains presented in the second row of Tab. 3 confirm that augmenting with data annotated optimally (i.e., using the combined real NuScenes train + Real val set) is beneficial. In particular, the performance of BEVFormer [14] and StreamPETR [34] improved significantly after the dataset was augmented with real validation set. The most notable gains for BEVFormer were observed in the mAP and NDS metrics, which increased by 3.23% and 2.92%, respectively. Similarly, for StreamPETR, increases in mAP and NDS were recorded at 3.08% and 2.02%, respectively.
Furthermore, augmentations using synthetic validation set yielded competitive improvements that almost matched the performance gains observed with real validation set. The gaps in performance metrics, such as mAP, NDS, and mAOE, were minimal, thus solidifying the value of synthetic augmentation compared to the train-only baseline. Specifically, BEVFormer [14] and StreamPETR [34] exhibited only slight gaps in mAP (0.54% and 1.85%), NDS (0.53% and 0.76%) and mAOE (2.17% and 1.72%), respectively. In addition, these discrepancies were even less pronounced compared to BEVControl*, highlighting the effectiveness of the Perl-based cross-attention mechanism.
4.3 Ablation Study
To determine the effectiveness of the fundamental components within our PerlDiff, we perform ablation studies concentrated on key elements: Perl-based cross-attention.
Effectiveness of Perl-based Cross-attention. To illustrate the impact of Perl-based cross-attention, we devised a comprehensive comparative experiment, the results of which are presented in Tab. 4. Method (a) employs road map and 3D box as conditions, which are integrated into the model using standard cross-attention [38, 6] with the configuration mirroring that of BEVControl*. “Box Mask” and “Road Mask” denote the process wherein the control information is merged with the model through Perl-based cross-attention. Method (a)$\rightarrow$(b) signifies the adoption of Perl-based cross-attention for the road map, leading to improvements of 0.44% in Road mIoU and 0.32% in NDS. These gains underscore the augmented controllability achieved by combining the road map with Perl-based cross-attention and its efficacy in aligning generated data with real observations. Additionally, Method (a)$\rightarrow$(c) results in marked improvements of 9.59% in mAP, 7.99% in NDS, 0.14% in mAOE and 3.80% in Vehicle mIoU, strongly supporting the utility of Perl-based cross-attention in producing accurate data-annotation alignments for objects. To optimally regulate elements of the background and foreground, Method (c), in contrast to baseline Method (a), indicates increases of 8.62% in mAP, 8.16% in NDS, 0.16% in mAOE, 0.52% in Road mIoU and 4.66% in Vehicle mIoU. These results validate the efficiency of Perl-based cross-attention in enhancing image controllability.
Table 4: Ablation of the Perl-based cross-attention mechanism, reporting 3D object detection improvements using BEVFormer [14] and BEV segmentation enhancements using CVT [41]. Numbers in parentheses indicate performance gains over the baseline.
@@ 4.3 Ablation Study
(a) 23.89 16.48 28.08 0.88 60.74 22.47
(b) 23.83 16.27 28.40 0.86 61.18 23.14
(c) 24.47 26.07 36.07 0.74 61.21 26.27
(d) 25.06 25.10 (+8.62%) 36.24 (+8.16%) 0.72 (-0.16%) 61.26 (+0.52%) 27.13 (+4.66%)
Table 5: Ablation of different values of masking map weight coefficients $\lambda_{s}$ and $\lambda_{b}$. We report the 3D object detection results based on BEVFormer [14] and BEV Segmentation results based on CVT [41].
@@
(a) 1.0 1.0 22.37 22.30 34.08 0.73 61.31 25.03
(b) 3.0 3.0 24.88 24.41 35.75 0.74 60.58 26.82
(c) 5.0 5.0 25.06 25.10 36.24 0.72 61.26 27.13
Effectiveness of Masking Map Weight Coefficients. Tab. 5 examines the effects of varying masking map weight coefficients $\lambda_{s}$ and $\lambda_{b}$, where higher weights indicate greater integration of Perl knowledge into network learning. The table demonstrates that the detection metrics improve with increasing values of $\lambda_{s}$ and $\lambda_{b}$, while the FID score increases, underscoring the significant role of Perl knowledge in the controllable learning process of the diffusion model. For optimal controllability of the generative model, we set the default values of $\lambda_{s}$ and $\lambda_{b}$ to 5.0 in the main paper.
Qualitative Visualization. We assess the efficacy of our PerlDiff against the baseline BEVControl*, with both approaches aimed at synthesizing perspective scene images. As illustrated in Fig. 1, PerlDiff generates images of substantially superior quality compared to the baseline, especially in the accurate depiction of object orientations, positions, and sizes. More results can be found in the Appendix D.
@@ 5 Conclusion
In conclusion, our PerlDiff introduces a streamlined framework that adeptly merges geometric constraints with synthetic street view image generation, harnessing diffusion models’ power for high-fidelity visuals. The architecture boasts a Perl-based controlling module (Perl-CM) that, through training, becomes seamlessly integrated with Stable Diffusion. Meanwhile, a cutting-edge Perl-based cross-attention mechanism guarantees meticulous feature guidance at the object level for precise control. Experiments on NuScenes [1] and KITTI [7] datasets confirm our PerlDiff’s enhanced performance in image synthesis and downstream tasks like 3D object detection and segmentation. Flexible yet precise, our PerlDiff’s method of Perl-based cross-attention with geometric perspective projections during training finely balances image realism with accurate condition alignment.
Limitation and Future Work. Fig. 6 depicts several failure cases of our PerlDiff, where the model erroneously generates vehicles with the front and rear orientations reversed, in contrast to the ground truth. This limitation arises from the usage of a Perl mask in our PerlDiff, which does not account for the orientation on the 2D Perl plane. Future endeavors may explore video generation, extending to work such as DrivingDiffusion [12], Panacea [36], and Driving into the Future [35].
Figure 6: Failure cases of our PerlDiff, with red markers highlighting instances where, compared to the ground truth, our PerlDiff generates images with the front and rear of vehicles reversed.
Appendix A DDPM Preliminaries

### v4

@@ 4.2 Main Results
Oracle BEVFormer 27.06 41.89 0.54
Oracle BEVFusion 35.54 41.20 0.56
MagicDrive [6] BEVFusion 16.20 12.30 23.32
MagicDrive* BEVFusion 15.92 10.27 20.42 0.78
BEVControl* BEVFusion 13.05 9.98 19.61 0.94
PerLDiff (Ours) BEVFusion 13.36 15.24 24.05 0.78
BEVGen [30] 25.54
BEVControl [38] BEVFormer 24.85 19.64 28.68 0.78
BEVControl* BEVFormer 13.05 16.48 28.08 0.88
MagicDrive* BEVFormer 15.92 15.21 28.79 0.81
PerLDiff (Ours) BEVFormer 13.36 25.10 36.24 0.72
Table 1: Comparison of the controllability on NuScenes validation set. BEVControl*, serving as the baseline, employs standard cross-attention mechanisms contrary to PerL-based cross-attention utilized in PerLDiff. MagicDrive* represents our replication using the official configuration. Bold indicates the best result.
4.2.1 Controllable Generation on NuScenes
To evaluate the effectiveness of PerLDiff, we trained the model on the NuScenes train set and subsequently generated a synthetic validation set using the provided road maps and 3D annotations. The controllability of PerLDiff was examined by applying perception models, originally trained on the real train set, to our synthetic validation set. As summarized in Tab. 1, PerLDiff outperforms competing methods across most metrics, as tested with BEVFormer [14] and BEVFusion [16]. In a rigorous comparison, we replicated BEVControl [38] using identical settings, with the exception of our innovative element: the PerL-based cross-attention mechanism. PerLDiff demonstrates notable improvements, with increases of 8.62%, 8.16%, and 0.16% in mean Average Precision (mAP), NuScenes Detection Score (NDS), and mean Average Orientation Error (mAOE), respectively, compared to BEVControl* when using BEVFormer. With BEVFusion, it achieves gains of 5.26%, 4.44%, and 0.16% in these metrics against BEVControl*, confirming its effectiveness at a resolution of $256\times 384$.
Method Oracle BEVGen BEVControl BEVControl* MagicDrive* PerLDiff (Ours)
@@ 4.2.1 Controllable Generation on NuScenes
Table 2: Comparison of the controllability performance on the NuScenes validation set using BEV segmentation metrics in CVT [41]. Bold indicates the best result.
@@
Table 3: Controllability comparison on KITTI [7] validation set, showcasing vehicle mAP obtained by MonoFlex [40] using data generated by PerLDiff and the baseline BEVControl*. “NuScenes $\rightarrow$ KITTI” denotes initial training on NuScenes train set followed by fine-tuning on KITTI train set. Bold indicates the best result.
@@
Table 4: Performance comparison for the boosting performance of 3D detection models using synthesized dataset on NuScenes [1] test set using BEVFormer [14] and StreamPETR [34]. The “train + Real val” configuration serves as a benchmark, representing the ideal upper performance limit achievable. “Syn. val*” represents the synthetic validation set generated by BEVControl. The numbers in parentheses indicate the performance disparity relative to the “train + Real val” configuration. Bold indicates the best result.
The superiority of PerLDiff is further affirmed by BEV segmentation metrics [41] presented in Tab. 2. It significantly outperforms BEVControl* with a 0.52% increase in Road mIoU and a 4.66% increase in Vehicle mIoU, validating the efficacy of the PerL-based cross-attention mechanism in enhancing scene controllability. Regarding the Fréchet Inception Distance (FID) [8] metric, our results are comparable to those of BEVControl*. While PerLDiff incorporates prior constraints to ensure accuracy in object detection, this may adversely affect the details in the background of the images. As illustrated in the Appendix, PerLDiff produces background details that do not align with those of real images.
Compared to the state-of-the-art MagicDrive [6], our method demonstrates superior performance across all metrics, particularly in the FID metric, reflecting an improvement of 2.84%. Additionally, we achieve a 2.94% improvement in mAP and a 0.73% increase in NDS. These results substantiate the strengths of PerLDiff in terms of both generation quality and controllability.
4.2.2 Controllable Generation on KITTI
The scarcity of training data in the KITTI dataset [7] often limits a generative model’s ability to understand the relationship between control information and image synthesis. To address this challenge, we implement two distinct strategies for generating images within the KITTI framework: one strategy involves direct training using the KITTI train set, while the other entails initial training on the NuScenes train set followed by fine-tuning on the KITTI train set. In Tab. 3, we present the results of monocular 3D object detection on the KITTI validation set, utilizing a pretrained MonoFlex [40] detector. As illustrated in Fig. 4, the naive approach results in significant misalignment between the labels and the corresponding generated images, leading to a considerable performance gap: 11.04 vs. 0.33 for one metric and 13.12 vs. 1.32 for another. There are two main reasons for the observed differences. First, the limited size of the KITTI dataset, which contains just 3,712 training images, impedes the learning process of traditional methods that do not utilize PerL masking map. Second, monocular 3D object detection is highly sensitive to accurate depth prediction. Depth is derived from the 2D projected size and the estimated 3D size through perspective projection. Our method produces more precise object sizing, thereby enhancing detection performance. More visual results on KITTI can be found in the Appendix.
@@ 4.2.2 Controllable Generation on KITTI
Figure 4: Qualitative visualization comparison on KITTI [7] dataset. Yellow markers denote instances where BEVControl* inaccurately generates output compared to PerLDiff.
@@
(a) 13.05 16.48 28.08 0.88 60.74 22.47
(b) 13.20 16.27 28.40 0.86 61.18 23.14
(c) 13.54 26.07 36.07 0.74 61.21 26.27
(d) 13.36 25.10 (+8.62%) 36.24 (+8.16%) 0.72 (-0.16%) 61.26 (+0.52%) 27.13 (+4.66%)
Table 5: Ablation of the PerL-based cross-attention, reporting 3D object detection improvements using BEVFormer and BEV segmentation enhancements using CVT. Numbers in parentheses indicate performance gains over the baseline. Bold indicates the best result.
4.2.3 Boosting Perception Models Using Synthesized Dataset
Generative models have become widely acknowledged as effective tools for data augmentation, thereby improving the generalization capabilities of perception models. To evaluate this, we leverage synthesized image dataset to improve the performance of various detection models on NuScenes test set. The gains presented in the second row of Tab. 4 confirm that augmenting with data annotated optimally (i.e., using the combined real NuScenes train + Real val set) is beneficial. The performance of BEVFormer [14] and StreamPETR [34] improved significantly after the dataset was augmented with real validation set. The most notable gains for BEVFormer were observed in the mAP and NDS metrics, which increased by 3.23% and 2.92%, respectively. Similarly, for StreamPETR, increases in mAP and NDS were recorded at 3.08% and 2.02%, respectively.
Furthermore, augmentations using synthetic validation set yielded competitive improvements that almost matched the performance gains observed with real validation set. The gaps in performance metrics, such as mAP, NDS, and mAOE, were minimal, thus solidifying the value of synthetic augmentation compared to the train-only baseline. Specifically, BEVFormer and StreamPETR exhibited only slight gaps in mAP (0.54% and 1.85%), NDS (0.53% and 0.76%) and mAOE (2.17% and 1.72%), respectively. In addition, these discrepancies were even less pronounced compared to BEVControl*, highlighting the effectiveness of the PerL-based cross-attention mechanism.
@@ 4.2.3 Boosting Perception Models Using Synthesized Dataset
(a) 1.0 1.0 12.87 22.30 34.08 0.73 61.31 25.03
(b) 3.0 3.0 14.03 24.41 35.75 0.74 60.58 26.82
(c) 5.0 5.0 13.36 25.10 36.24 0.72 61.26 27.13
(d) 10.0 10.0 14.24 24.98 35.52 0.76 61.75 26.62
Table 6: Ablation of different values of masking map weight coefficients $\lambda_{s}$ and $\lambda_{b}$ in PerL-based cross-attention. We report the 3D object detection results based on BEVFormer and BEV Segmentation results based on CVT. Bold indicates the best result.
4.3 Ablation Study
To determine the effectiveness of the key components within PerLDiff, we perform ablation studies concentrated on key elements: PerL-based cross-attention. Additionally, we provide more qualitative results in the Appendix to further demonstrate the efficacy of PerLDiff compared to the baseline approaches, BEVControl* and MagicDrive.
4.3.1 Effectiveness of PerL-based Cross-attention
To illustrate the impact of PerL-based cross-attention, we devised a comprehensive comparative experiment, the results of which are presented in Tab. 5. Method (a) employs road map and 3D box as conditions, which are integrated into the model using standard cross-attention [38, 6] with the configuration mirroring that of BEVControl*. “Box Mask” and “Road Mask” denote the process wherein the control information is merged with the model through PerL-based cross-attention. Method (a)$\rightarrow$(b) signifies the adoption of PerL-based cross-attention for the road map, leading to improvements of 0.44% in Road mIoU and 0.32% in NDS. These gains underscore the augmented controllability achieved by combining the road map with PerL-based cross-attention and its efficacy in aligning generated data with real observations. Additionally, Method (a)$\rightarrow$(c) results in marked improvements of 9.59% in mAP, 7.99% in NDS, 0.14% in mAOE and 3.80% in Vehicle mIoU, strongly supporting the utility of PerL-based cross-attention in producing accurate data-annotation alignments for objects. To optimally regulate elements of the background and foreground, Method (c), in contrast to baseline Method (a), indicates increases of 8.62% in mAP, 8.16% in NDS, 0.16% in mAOE, 0.52% in Road mIoU and 4.66% in Vehicle mIoU. These results validate the efficiency of PerL-based cross-attention in enhancing image controllability.
4.3.2 Effectiveness of Masking Map Weight Coefficients
Tab. 6 examines the effects of varying the masking map weight coefficients $\lambda_{s}$ and $\lambda_{b}$, where higher values indicate a greater integration of PerL knowledge into network learning. The table demonstrates that detection metrics improve with increasing values of $\lambda_{s}$ and $\lambda_{b}$ within a certain range. However, the FID score also increases, underscoring the significant role of PerL knowledge in the controllable learning process. For optimal controllability, we set the default values of $\lambda_{s}$ and $\lambda_{b}$ to 5.0 in the main text.
4.3.3 Effectiveness of Different Road Map Encoders
In Tab. 7, we present an ablation study comparing different road map encoders, replacing ConvNext [15] with alternatives such as the CLIP image encoder. ConvNext was initially chosen for its ability to effectively extract image features while maintaining a relatively low parameter count (27.82M). Additionally, it has demonstrated strong performance in downstream segmentation tasks, particularly in capturing edge information [15]. As shown in Tab. 7, the results indicate that the choice of feature extraction network has minimal impact on PerLDiff, with only a ±$\pm 0.02$ NDS difference between the CLIP and ConvNext encoders. Therefore, we adopt ConvNext as the road map encoder and freeze its weights in the main results.
[…]
CLIP encoder BEVFormer 303.97 12.77 25.41 36.26 0.73
ConvNext encoder 27.82 13.36 25.10 36.24 0.72
CLIP encoder BEVFusion 303.97 12.77 15.80 24.45 0.81
ConvNext encoder 27.82 13.36 15.24 24.05 0.78
Table 7: Ablation study comparing different road map encoders with frozen weights. We present 3D object detection results based on BEVFormer and BEVFusion. Bold indicates the best result.
@@ 5 Conclusion
In conclusion, PerLDiff introduces a streamlined framework that adeptly merges geometric constraints with synthetic street view image generation, harnessing diffusion models’ power for high-fidelity visuals. The architecture boasts a PerL-based controlling module (PerL-CM), through training, becomes seamlessly integrated with Stable Diffusion. Meanwhile, a cutting-edge PerL-based cross-attention mechanism guarantees meticulous feature guidance at the object level for precise control. Experiments on NuScenes and KITTI datasets confirm PerLDiff’s enhanced performance in image synthesis and downstream tasks like 3D object detection and segmentation. Flexible yet precise, PerLDiff’s method of PerL-based cross-attention with geometric perspective projections during training finely balances image realism with accurate condition alignment.
Acknowledgement
This work was supported by National Natural Science Foundation of China (No. 62476051, No. 62176047) and Sichuan Natural Science Foundation (No. 2024NSFTD0041).
Supplementary Material
The supplementary material is organized into the following sections:
  • Section A: DDPM Preliminaries
  • Section B: Implementation Details
  • Section C: Limitation and Future Work
  • Section D: Additional Experiments
  • Section E: Visualization Results
A DDPM Preliminaries